# Indirect Influence on Network Diffusion

### v1

@@ Introduction
One proposed approach to incorporate these indirect connections is through the use of the $d$-path Laplacian and the $d$-path adjacency matrices 13, a natural generalization of the Laplacian and the adjacency matrices of direct connections. This framework has been successfully applied in various contexts, such as capturing the indirect diffusion of innovations in social networks 14 or in modeling epidemic spreading with long-range diffusion 15.
Although the practical use of this framework has been extensively considered in the literature 13, 14, 15, 16, little is known about the interplay of the network structure and the diffusion dynamics, which has proven to be key in other frameworks 17, 18.
@@
To analyze diffusion dynamics incorporating non-local interactions, we employ the $d$-path Laplacian framework 13. This approach extends the standard Laplacian model by considering pathways of length $d>1$ between nodes. The framework uses $d$-path adjacency matrices, $A^{(d)}$, where $\left(A^{(d)}\right)_{ij}=a^{(d)}_{ij}=1$ if the shortest path distance between nodes $i$ and $j$ is exactly $d$, and 00 otherwise. From these, the corresponding $d$-path Laplacian, $L^{(d)}=K^{(d)}-A^{(d)}$, is constructed, where $K^{(d)}$ is the diagonal matrix of $d$-path degrees $k^{(d)}_{i}=\sum_{j}a^{(d)}_{ij}$ (see Methods).
@@
where $0=\hat{\lambda}_{1}<\hat{\lambda}_{2}\leq\cdots\leq\hat{\lambda}_{N}$ are the eigenvalues and $\hat{\mathbf{v}}_{i}$ are the corresponding eigenvectors of $\hat{L}$, and $C_{i}$ are constants determined by initial conditions. The convergence to the equilibrium state, where $x_{i}=x_{j}$ for all $i$, $j$, is governed by the second smallest eigenvalue, $\hat{\lambda}_{2}$, often called the algebraic connectivity or Fiddler value 19, 20. The characteristic diffusion time is $\tau=1/\hat{\lambda}_{2}$; a larger $\hat{\lambda}_{2}$ means faster convergence.
In contrast, standard diffusion models only consider direct connections, described by $\dot{\mathbf{x}}=L\mathbf{x}$, where $L=L^{(1)}$, with a diffusion time $\tau=1/\lambda_{2}$, where $\lambda_{2}$ is the second smallest eigenvalue of the standard Laplacian $L$. To quantify the impact of including indirect pathways on the diffusion timescale, we introduce the indirect influence, defined as the normalized change in the algebraic connectivity:
@@
This parameter ranges from 0 to 1, since $N$ is the maximum value that any eigenvalue can take in a Laplacian matrix 19, and $\hat{\lambda}_{2}\geqslant\lambda_{2}$ since the original 1-path network is a (weighted) spanning graph of the path network 21. A value of $\zeta=0$ indicates that indirect connections do not alter the diffusion timescale compared to the direct-only model, while $\zeta>0$ means that indirect pathways accelerate the diffusion process, with larger values indicating a greater impact. This parameter provides a quantitative measure to address when and how significantly indirect connections reshape network diffusion dynamics.
@@
In the large $\alpha$ limit, the path Laplacian $\hat{L}$ is dominated by the $d=1$ term. We can approximate it by including only the first significant indirect contribution, the $d=2$ term, scaled by a small parameter $\epsilon=2^{-\alpha}\ll 1$:
$\hat{L}\approx L+2^{-\alpha}L^{(2)}=L+\epsilon L^{(2)}.$ (5)
@@
$\hat{\lambda}_{2}\approx\lambda_{2}+\epsilon\mathbf{v}_{2}^{T}L^{(2)}\mathbf{v}_{2}=\lambda_{2}+2^{-\alpha}\mathbf{v}_{2}^{T}L^{(2)}\mathbf{v}_{2},$ (6)
@@
$\zeta\approx\frac{2^{-\alpha}\mathbf{v}_{2}^{T}L^{(2)}\mathbf{v}_{2}}{N}.$ (7)
This expression quantifies the indirect influence arising from $2$-paths in the perturbative regime, but it still depends on the specific eigenvector $\mathbf{v}_{2}$ and the $2$-path Laplacian $L^{(2)}$ of a given network.
To obtain an expression dependent only on average network properties, we employ a mean-field approach to approximate the expectation value of the quadratic form $\mathbf{v}_{2}^{T}L^{(2)}\mathbf{v}_{2}$ over an ensemble of general random networks. Assuming the entries of the eigenvector $\mathbf{v}_{2}$ are uncorrelated with the entries of $L^{(2)}$ and are independent identically distributed random variables satisfying $\sum_{i=1}^{N}(\mathbf{v}_{2})_{i}^{2}=1$ and $\sum_{i=1}^{N}(\mathbf{v}_{2})_{i}=0$ 19, we can approximate the expectation:
@@
Defining the expected connection probability of the $2$-path network as $p^{(2)}\equiv\big{\langle}a^{(2)}_{ij}\big{\rangle}$, and noting that the expected $2$-path degree is $\big{\langle}k^{(2)}\big{\rangle}\approx(N-1)p^{(2)}$, we arrive at:
@@
The expected $2$-path connection probability $p^{(2)}$ can be expressed in terms of the direct expected connection probability $p$ of the underlying network. A $2$-path exists between nodes $i$ and $j$ if they are not directly connected ($a_{ij}=0$) but there exists at least one intermediate node $k$ such that $a_{ik}=1$ and $a_{kj}=1$. Using a mean-field argument for the existence probability over the $N-2$ potential intermediate nodes:
@@
Figure 2: Expected connection probabilities of $\bm{d}$-path networks relative to the 1-path network. Expected connection probabilities for different $d$-path networks as a function of the expected 1-path connection probability, with ${N=1000}$. a Erdős-Rényi (ER) networks and b Random Regular (RR) networks. Each point corresponds to the average over 50 different networks with error bars of one standard deviation, and the solid lines correspond to their respective analytical expression. The gray shaded range indicates the values of ${p<\log N/N}$, corresponding to the values where the network is almost surely disconnected 22, and therefore are not considered.
@@
Figure 3: Average indirect influence across parameter space. Average indirect influence computed over all possible $d=1,\dots,D$ paths for networks with $N=1000$. a Erdős-Rényi (ER) networks. b Random Regular (RR) networks. Solid white lines correspond to the contour levels of the color map that corresponds to the simulations, and the dashed white lines correspond to the theoretical contour levels. We compute 100 different networks for each combination of $\alpha$ ranging from 0 to 4 in steps of 0.1 and $p$ ranging from 00 to $1$ in steps of 0.001.
@@
If we consider only direct and $2$-path connections, the path Laplacian is $\hat{L}=L^{(1)}+L^{(2)}$. This structure resembles the Laplacian used in studies of diffusion on duplex networks 17, 23. Analogously, we can approximate the combined $1$-path and $2$-path network as a single effective network. Since the edge sets of the $1$-path and $2$-path networks are disjoint, by definition of $A^{(2)}$, we approximate this effective network as an Erdős-Rényi network with $N$ nodes and an effective connection probability $p_{\text{eff}}=p+p^{(2)}$, where $p$ and $p^{(2)}$ are given by Eq. (12). The algebraic connectivity $\hat{\lambda}_{2}$ for this truncated $\alpha=0$ case can then be approximated by the known results for the algebraic connectivity of ER and RR networks 24, 25. The indirect influence is then approximated as:
$\zeta_{\alpha=0}^{\text{ER/RR}}\approx\frac{\lambda^{\text{ER}}_{2}(N,p+p^{(2)})-\lambda^{\text{ER/RR}}_{2}(N,p)}{N}.$ (15)
The approximation for large ($N\gg 1$) unweighted RR networks of degree $k=p(N-1)$ is:
@@
$\begin{split}\lambda^{\text{ER}}_{2}(N,p)\approx p(N-1)&-\sqrt{2p(1-p)(N-1)\log N}\\ &+\sqrt{\frac{(N-1)p(1-p)}{2\log N}}\log\sqrt{2\pi\log\left(\frac{N^{2}}{2\pi}\right)}\\ &-\sqrt{\frac{(N-1)p(1-p)}{2\log N}}\,\gamma\,,\end{split}$ (17)
@@
Figure 4: Degree distribution of the path network. Degree distribution of the path network, $k=k^{(1)}+k^{(2)}$, which includes 1-path and 2-path links, averaged over 50 realizations of a ER networks and b RR networks. In both cases, $N=1000$ and $\big{\langle}k^{(1)}\big{\rangle}=30$. Solid lines represent the theoretical binomial distribution that an ER network should have, with $N$ nodes and a connection probability of $p+p^{(2)}$.
@@
If we consider the full path Laplacian at $\alpha=0$, assuming the underlying network is connected, the path network effectively becomes the complete graph $K_{N}$. This is because for any two nodes $i$, $j$ in a connected graph, there exists a shortest path of some length $d\leq D$. The algebraic connectivity of the complete unweighted network is $\lambda_{2}=N$  19. Therefore, in this limit, the indirect influence reaches its maximum possible value:
$\zeta_{\alpha=0}^{\text{ER/RR}}\approx\frac{N-\lambda^{\text{ER/RR}}_{2}(N,p)}{N}.$ (18)
@@
Inspired by the concept of connectivity transitions in random graphs 22, we focus on the completeness of the network formed by considering both direct and indirect paths. For simplicity, let’s first consider the network formed by only $1$-paths and $2$-paths, ignoring the weighting $d^{-\alpha}$ and focusing solely on the presence or absence of connections. The density of connections in this combined network can be quantified by its completeness $\nu$, defined as the probability that a randomly chosen pair of nodes $i$ and $j$ is connected by either a 1-path or a 2-path:
@@
We consider this completeness $\nu$ as an order parameter, with the direct expected connection probability $p$ acting as the control parameter. We look for a phase transition, marked by a critical value $p_{c}(N)$, where the probability of the combined network being fully connected sharply transitions from 00 to $1$ as $N\rightarrow\infty$, i.e.,
@@
$\begin{split}P\{\nu=1\}&=1-P\{\exists i,j\,|\,a_{ij}=a^{(2)}_{ij}=0\}\\ &=1-\left(1-\left(p+p^{(2)}\right)\right)=p+p^{(2)}\\ &=p+(1-p)\left(1-\left(1-p^{2}\right)^{N-2}\right)\\ &=1-\left(1-p^{2}\right)^{N}\\ &\approx 1-e^{-p^{2}N}\end{split}$ (21)
@@ Methods
We model diffusion dynamics on a network represented as an unweighted, connected graph $G=(V,E)$, where $V$ is the set of $N$ nodes (indexed $i=1,...,N$) and $E$ is the set of edges representing direct connections between nodes. To incorporate the influence of interactions beyond direct neighbors, we employ the $d$-path Laplacian framework 13.
@@
[…]
@@
$\begin{split}p^{(d)}&=P\{a_{ij}=0,...,a^{(d-1)}_{ij}=0\}(1-P\{\forall k_{1},...,k_{d-1}\in G\,|\,a_{ik_{1}}...a_{k_{d-1}j}=0\})\\ &\approx P\{a_{ij}=0,...,a^{(d-1)}_{ij}=0\}\left(1-(1-P\{k_{1},...,k_{d-1}\in G\,|\,a_{ik_{1}}...a_{k_{d-1}j}=1\})^{\binom{N-2}{d-1}}\right).\end{split}$ (40)
@@
$p^{(d)}=\left(1-\sum_{i=1}^{d-1}p^{(i)}\right)\left(1-\left(1-p^{d}\right)^{\binom{N-2}{d-1}}\right).$ (41)

### v2

@@ Introduction
One proposed approach to incorporate these indirect connections is through the use of the $d$-path Laplacian and the $d$-path adjacency matrices 13, a natural generalization of the Laplacian and the adjacency matrices to include higher-order paths and establish the theoretical possibility of non-local diffusion, even demonstrating super-diffusive behavior under certain conditions for single-particle dynamics14, 15. Other options to incorporate indirect influence in networks include the multi-hop random walks model16, and the fractional Laplacian17, 18, which show important qualitative and quantitative differences with respect to the $d$-path Laplacian approach. Beyond diffusion, indirect influence can also affect other dynamics, such as synchronization19 and consensus models20.
Indirect diffusion has proven effective in modeling real-world scenarios, e.g., capturing the long-range dispersal of pathogens via vectors in epidemics 21, showing how peer pressure can shape consensus and leadership in social groups22, and accounting for the accelerated diffusion of innovations through socially-close peers in experiments 23. Despite its practical importance, the interplay between network structure and diffusion dynamics remains poorly understood, although it has proven central in other frameworks 24, 25.
@@ Results
To analyze diffusion dynamics incorporating non-local interactions, we employ the $d$-path Laplacian framework 13. This approach extends the standard Laplacian model by considering pathways of length $d>1$ between nodes. The framework uses $d$-path adjacency matrices, $A^{(d)}$, where $\left(A^{(d)}\right)_{ij}=a^{(d)}_{ij}=1$ if the shortest path distance between nodes $i$ and $j$ is exactly $d$, and 0 otherwise. From these, the corresponding $d$-path Laplacian, $L^{(d)}=K^{(d)}-A^{(d)}$, is constructed, where $K^{(d)}$ is the diagonal matrix of $d$-path degrees $k^{(d)}_{i}=\sum_{j}a^{(d)}_{ij}$ (see Methods).
@@
where $0=\hat{\lambda}_{1}<\hat{\lambda}_{2}\leq\cdots\leq\hat{\lambda}_{N}$ are the eigenvalues and $\hat{\mathbf{v}}_{i}$ are the corresponding eigenvectors of $\hat{L}$, and $C_{i}$ are constants determined by initial conditions. The convergence to the equilibrium state, where $x_{i}=x_{j}$ for all $i$, $j$, is governed by the second smallest eigenvalue, $\hat{\lambda}_{2}$, often called the algebraic connectivity or Fiedler value 26, 27. The characteristic diffusion time is $\tau=1/\hat{\lambda}_{2}$; a larger $\hat{\lambda}_{2}$ means faster convergence.
In contrast, standard diffusion models only consider direct connections, described by $\dot{\mathbf{x}}=-L\mathbf{x}$, where $L=L^{(1)}$, with a diffusion time $\tau=1/\lambda_{2}$, where $\lambda_{2}$ is the second smallest eigenvalue of the standard Laplacian $L$. To quantify the impact of including indirect pathways on the diffusion timescale, we introduce the indirect influence, defined as the normalized change in the algebraic connectivity:
@@
This parameter ranges from 0 to 1, since $N$ is the maximum value that any eigenvalue can take in a Laplacian matrix 26, and $\hat{\lambda}_{2}\geqslant\lambda_{2}$ since the original 1-path network is a (weighted) spanning graph of the path network 28. A value of $\zeta=0$ indicates that indirect connections do not alter the diffusion timescale compared to the direct-only model, while $\zeta>0$ means that indirect pathways accelerate the diffusion process, with larger values indicating a greater impact. This parameter provides a quantitative measure to address when and how significantly indirect connections reshape network diffusion dynamics.
@@
In the large $\alpha$ limit, the path Laplacian $\hat{L}$ is dominated by the $d=1$ term. We start by including the most significant indirect contribution, the $d=2$ term, scaled by a small parameter $\epsilon=2^{-\alpha}\ll 1$:
$\hat{L}=L+2^{-\alpha}L^{(2)}=L+\epsilon L^{(2)}.$ (5)
@@
$\hat{\lambda}_{2}=\lambda_{2}+\epsilon\mathbf{v}_{2}^{T}L^{(2)}\mathbf{v}_{2}+O(\epsilon^{2}).$ (6)
@@
$\zeta\approx\frac{2^{-\alpha}\mathbf{v}_{2}^{T}L^{(2)}\mathbf{v}_{2}}{N},$ (7)
to first order in $\epsilon$. This expression quantifies the indirect influence arising from $2$-paths in the perturbative regime, but it still depends on the specific eigenvector $\mathbf{v}_{2}$ and the $2$-path Laplacian $L^{(2)}$ of a given network.
To obtain an expression dependent only on average network properties, we employ a mean-field approach to approximate the expectation value of the quadratic form $\mathbf{v}_{2}^{T}L^{(2)}\mathbf{v}_{2}$ over an ensemble of general random networks. Assuming the entries of the eigenvector $\mathbf{v}_{2}$ are uncorrelated with the entries of $L^{(2)}$ and are independent identically distributed random variables satisfying $\sum_{i=1}^{N}(\mathbf{v}_{2})_{i}^{2}=1$ and $\sum_{i=1}^{N}(\mathbf{v}_{2})_{i}=0$ 26, we can approximate the expectation:
@@
Defining the expected connection probability of the $2$-path network as $p^{(2)}\equiv\big\langle a^{(2)}_{ij}\big\rangle$, and noting that the expected $2$-path degree is $\big\langle k^{(2)}\big\rangle=(N-1)p^{(2)}$, we arrive at:
@@
The expected $2$-path connection probability $p^{(2)}$ can be expressed in terms of the direct expected connection probability $p$ of the underlying network. A $2$-path exists between nodes $i$ and $j$ if they are not directly connected ($a_{ij}=0$) but there exists at least one intermediate node $k$ such that $a_{ik}=1$ and $a_{kj}=1$. Using a mean-field argument for the existence probability over the N2N-2 potential intermediate nodes:
@@
Figure 2: Expected connection probabilities of $\bm{d}$-path networks relative to the 1-path network. Expected connection probabilities for different $d$-path networks as a function of the expected 1-path connection probability, with N=1000{N=1000}. a Erdős-Rényi (ER) networks and b Random Regular (RR) networks. Each point corresponds to the average over 50 different networks with error bars of one standard deviation, and the solid lines correspond to their respective analytical expression. The gray shaded range indicates the values of $p<\log N/N$}, corresponding to the values where the network is almost surely disconnected 29, and therefore are not considered.
@@
Figure 3: Average indirect influence across parameter space. Average indirect influence computed over all possible $d=1,\dots,D$ paths for networks with $N=1000$. a Erdős-Rényi (ER) networks. b Random Regular (RR) networks. Solid white lines correspond to the contour levels of the color map that corresponds to the simulations, and the dashed white lines correspond to the theoretical contour levels. We compute 100 different networks for each combination of $\alpha$ ranging from 0 to 4 in steps of 0.1 and $p$ ranging from 0 to 11 in steps of 0.001.
@@
If we consider only direct and $2$-path connections, the path Laplacian is $\hat{L}=L^{(1)}+L^{(2)}$. This structure resembles the Laplacian used in studies of diffusion on duplex networks 24, 30. Analogously, we can approximate the combined $1$-path and $2$-path network as a single effective network. Since the edge sets of the $1$-path and $2$-path networks are disjoint, by definition of $A^{(2)}$, we approximate this effective network as an Erdős-Rényi network with $N$ nodes and an effective connection probability $p_{\text{eff}}=p+p^{(2)}$, where $p$ and $p^{(2)}$ are given by Eq. (12). The algebraic connectivity $\hat{\lambda}_{2}$ for this truncated $\alpha=0$ case can then be approximated by the known results for the algebraic connectivity of ER and RR networks 31, 32. The indirect influence can be written as:
$\zeta_{\alpha=0}^{\text{ER/RR}}=\frac{\lambda^{\text{ER}}_{2}(N,p+p^{(2)})-\lambda^{\text{ER/RR}}_{2}(N,p)}{N}.$ (15)
The approximation for large ($N\gg 1$) unweighted RR networks of degree k=p(N1)k=p(N-1) is:
@@
[…]
@@
Figure 4: Degree distribution of the path network. Degree distribution of the path network, $k=k^{(1)}+k^{(2)}$, which includes 1-path and 2-path links, averaged over 50 realizations of a ER networks and b $R$ networks. In both cases, $N=1000$ and $\big\langle k^{(1)}\big\rangle=30$. Solid lines represent the theoretical binomial distribution that an ER network should have, with $N$ nodes and a connection probability of $p+p^{(2)}$.
@@
If we consider the full path Laplacian at $\alpha=0$, assuming the underlying network is connected, the path network effectively becomes the complete graph $K_{N}$. This is because for any two nodes $i$, $j$ in a connected graph, there exists a shortest path of some length $d\leq D$. The algebraic connectivity of the complete unweighted network is $\lambda_{2}=N$ 26. Therefore, in this limit, the indirect influence reaches its maximum possible value:
$\zeta_{\alpha=0}^{\text{ER/RR}}=\frac{N-\lambda^{\text{ER/RR}}_{2}(N,p)}{N}.$ (18)
@@
Inspired by the concept of connectivity transitions in random graphs 29, we focus on the completeness of the network formed by considering both direct and indirect paths. For simplicity, let’s first consider the network formed by only $1$-paths and $2$-paths, ignoring the weighting $d^{-\alpha}$ and focusing solely on the presence or absence of connections. The density of connections in this combined network can be quantified by its completeness $\nu$, defined as the probability that a randomly chosen pair of nodes $i$ and $j$ is connected by either a 1-path or a 2-path:
@@
We consider this completeness $\nu$ as an order parameter, with the direct expected connection probability $p$ acting as the control parameter. We look for a phase transition, marked by a critical value $p_{c}(N)$, where the probability of the combined network being fully connected sharply transitions from 0 to $1$ as $N\rightarrow\infty$, i.e.,
@@
$\begin{split}P\{\nu=1\}&=1-P\{\exists i,j\,|\,a_{ij}=a^{(2)}_{ij}=0\}\\ &=1-\left(1-\left(p+p^{(2)}\right)\right)=p+p^{(2)}\\ &=p+(1-p)\left(1-\left(1-p^{2}\right)^{N-2}\right)\\ &=1-\left(1-p^{2}\right)^{N}\\ &=1-e^{-p^{2}N}\end{split}$ (21)
@@
Discussion
This work provides a quantitative framework for analyzing the impact of indirect pathways on network diffusion. The central results, including the indirect influence parameter $\zeta$, its predictive phase diagram, and the critical threshold for the onset of $d$-paths, $p_{c}(N)$, offer a practical method for assessing the significance of non-local interactions. The theory, derived for analytically tractable random networks, is validated by numerical simulations, confirming that indirect pathways emerge as a collective structural phenomenon rather than a gradual effect.
The theoretical approach relies on mean-field approximations that are well-suited for random graphs. This scope invites future work on networks with more complex topologies, such as those with high clustering or community structure, where structural correlations would likely alter the quantitative predictions. A key implication of our framework is the potential for empirical measurement of $\zeta$. Observing the global diffusion timescale of a process on a network with a known structure would allow for the calculation of an empirical $\zeta$. This value could then be used to test the model and estimate system-specific parameters, such as the influence decay rate $\alpha$, bridging the gap between theory and real-world data.
The model employs a power-law decay of influence $d^{-\alpha}$, a flexible choice corresponding to a Mellin transform. While other functional forms, like exponential decay, are possible, the qualitative features of the phase transition are expected to be robust. In conclusion, this study’s findings provide a rigorous guide for model selection in network science, with direct applications in fields such as epidemiology and the study of social influence.
@@ Methods
We model diffusion dynamics on a network represented as an unweighted, connected graph $G=(V,E)$, where $V$ is the set of $N$ nodes (indexed i=1,,Ni=1,...,N) and $E$ is the set of edges representing direct connections between nodes. To incorporate the influence of interactions beyond direct neighbors, we employ the $d$-path Laplacian framework 13.
@@
$\sum_{i=1}^{N}(\mathbf{v}_{2})_{i}=0\ \Rightarrow\ 0=\left(\sum_{i=1}^{N}(\mathbf{v}_{2})_{i}\right)\left(\sum_{j=1}^{N}(\mathbf{v}_{2})_{j}\right).$ (36)
@@
[…]